OpenStudy (anonymous):

What is the solution for this system of linear equations: Check my answer please, y=-4x+23 y=7x-43 so 7x-43=-4x+23 subtracted 7 from both sides -43=-11+23 then subtract 23 from both sides -66 divided by -11 = 6 so x=6 ? :)

OpenStudy (anonymous):

can you help me find y?

jimthompson5910 (jim_thompson5910):

x = 6 is correct

jimthompson5910 (jim_thompson5910):

nice work so far

jimthompson5910 (jim_thompson5910):

pick any equation with x and y in it say you pick the first one

jimthompson5910 (jim_thompson5910):

y = -4x + 23 now plug in x = 6 and evaluate to find y y = -4(6) + 23 y = -24 + 23 y = -1

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

you can also use the second equation to get the same y value y=7x-43 y=7(6)-43 ... plug in x = 6 y = 42 - 43 y = -1

jimthompson5910 (jim_thompson5910):

So that's a good way to check. If you get two different y values for some given x value, then something went wrong.

OpenStudy (anonymous):

Thank you so much! I was confused on finding the y. I was just putting x in the place of the x in the original question. Thank you!! :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

remember set them side by side for this one subtract 23 on both sides and subtract 7 from both sides it should look like this -4x+23=7x-43 -7 -23 -7 -23 ------------- -11x=-66 x=6 and now plug in 6 in any of the first equations Lets pick the first one to get y -4(6)+23 -24+23 y=-1